# arXiv:2603.00942v1[math.DG] 01 Mar 2026

### v1

@@ 1 Introduction
In the past four decades, smooth metric measure spaces have been extensively investigated in geometric analysis, and have close relation with probability theory and optimal transport. We wish to mention again several important examples listed in our previous work [67, Section 1] to support this viewpoint — G. Perelman’s pioneering work [57] on the entropy formula of the Ricci flow under a weighted measure, Munteanu-Wang’s series works [54, 55, 56] about geometry of manifolds with weighted densities, Wei-Wylie’s interesting work [63] on some comparison theorems for manifolds with Bakry-Émery Ricci curvature bounded. Besides, recently, by mainly using the rearrangement technique and suitably constructing trial functions, under the constraint of fixed weighted volume, R. F. Chen and J. Mao [20] successfully obtain several isoperimetric inequalities for the first and the second Dirichlet eigenvalues, the first nonzero Neumann eigenvalue of the Witten Laplacian on bounded domains in space forms. These spectral isoperimetric inequalities extend the classical ones (i.e., the Faber-Krahn inequality, the Hong-Krahn-Szegő inequality, and the Szegő-Weinberger inequality) of the Laplacian.
@@
$\lim\limits_{t\rightarrow 0}\int\limits_{M^{n}}f(x)\delta^{\phi}_{y}(x)e^{-\phi}dv(x)=f(y).$
@@
For detailed calculations, see [65, Appendix].
@@
There are so many mathematicians having interest on the (non-weighted or weighted) heat kernel and its applications, and many interesting results have been obtained. Here we wish to give a partial list of those results which influence me a lot. In 1981, S. Y. Cheng, P. Li and S. T. Yau [25] gave uniform Gaussian estimates for the heat kernel on Riemannian manifolds with sectional curvature bounded from below. Their estimates were improved by J. Cheeger, M. Gromov, M. Taylor [16] to manifolds with bounded geometry. By using the gradient estimate technique and the Harnack inequality, P. Li and S. T. Yau [42] obtained sharp Gaussian upper and lower bounds for the Schrödinger operator on Riemannian manifolds with nonnegative Ricci curvature. For smooth metric measure spaces with $\mathrm{Ric}^{N}_{\phi}$ ($N<\infty$) bounded from below by a negative quadratic function, X. D. Li [41] got Gaussian estimates for the weighted heat kernel, and proved an $L^{1}_{\phi}$-Liouville theorem. In [64], J. Y. Wu obtained weighted heat kernel estimates on smooth metric measure spaces with the weighted Ricci curvature bounded from below by a negative constant and the potential function bounded. By using the De Giorgi-Nash-Moser theory and the weighted version of Davies’s integral estimate [26], J. Y. Wu and P. Wu [65, Theorem 1.1] successfully gave a local Gaussian upper bound for the weighted heat kernel on complete noncompact smooth metric measure spaces $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$ with $\mathrm{Ric}_{\phi}\geq 0$. Their upper bound estimate generalizes the classical result of Li-Yau [42]. As applications of their local Gaussian upper bound estimate for the weighted heat kernel, they showed that if furthermore the potential function $\phi$ has at most quadratic growth, then a sharp $L^{1}_{\phi}$-Liouville theorem for $\phi$-subharmonic functions and an $L^{1}_{\phi}$-uniqueness property for nonnegative solutions of the weighted heat equation can be obtained (see [65, Theorem 1.5]). Besides, it is not hard to see from [13, 38] that any complete noncompact shrinking or steady gradient Ricci soliton has nonnegative weighted Ricci curvature and its potential function $\phi$ should have at most quadratic growth. Then together with Wu-Wu’s $L^{1}_{\phi}$-Liouville theorem mentioned above, it is easy to know that any $L^{1}_{\phi}$-integral $\phi$-subharmonic function on gradient shrinking and steady Ricci solitons must be constant (see [65, Corollary 1.6]). Later, as a sequel of the work [65], by mainly using the De Giorgi-Nash-Moser theory, J. Y. Wu and P. Wu [66, Theorem 1.1] obtained local Gaussian (upper and lower) bounds for the weighted heat kernel on complete noncompact smooth metric measure spaces $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$ with Bakry-Émery Ricci curvature bounded from below by some nonnegative constant (i.e. $\mathrm{Ric}_{\phi}\geq-(n-1)K$ for some constant $K\geq 0$). Moreover, when the curvature assumption was reduced to $\mathrm{Ric}_{\phi}\geq 0$ (i.e. $K=0$), they pointed out that the local Gaussian lower bound for the weighted heat kernel can be achieved by the $\phi$-heat kernel for steady Gaussian soliton (explicitly given in Example 1.3) as long as the time $t$ is large enough. Besides, as applications, they used their local Gaussian (upper and lower) bounds for the weighted heat kernel to obtain estimates for closed eigenvalues and for $\phi$-Green’s function of the Witten-Laplacian – see [66, Sections 6 and 7] for details. In our recent work [67], by utilizing some functional inequalities and combining with De Giorgi-Moser iteration technique, some estimates for the weighted $L^{\infty}$-norm of eigenfunctions of the Witten-Laplacian on smooth metric measure spaces have been given. Besides, we also proved the regularity and uniqueness of the weighted heat kernel on smooth metric measure spaces, which generalized Chavel-Feldman’s related results in [14].
@@
In this paper, we wish to give a comparison theorem for the weighted heat kernel (of the weighted heat equation) on complete manifolds with radial curvatures bounded. In this setting, the model space would be spherically symmetric manifolds. In order to state our main result clearly, we need to recall some preliminary knowledge about the exponential mapping, the cut locus, the radial direction, the injectivity radius, the radial (Ricci or sectional) curvature having a bound given by a continuous function (see Definitions 2.2 and 2.3), spherically symmetric manifolds (see Definition 2.1), and so on. However, in order to control the length of this section, we wish to give preliminaries in Section 2 of this paper. By the way, these preliminaries can also be found in our previous works [31, Section 1], [46, Section 2.1 of Chapter 2], [47, Section 2].
@@
which will be used to determine our model space $0,l)\times_{f}\mathbb{S}^{n-1}$ (i.e. spherically symmetric manifolds, which are called generalized space forms in [40]). See Definition 2.1 in Section 2 for the precise statement such that a domain to be spherically symmetric. Readers can check our previous work [31, pp. 705–709] for a brief introduction and an interesting spectral asymptotic property of spherically symmetric manifolds. Obviously, if in (1.5), $\kappa(s)\equiv K$ for some constant $K$, then
@@
(see [36, Subsection 4.6]) Let $\mathbb{H}^{3}$ be the hyperbolic $3$-space with constant666 For the weighted heat kernel $H^{\phi}$ shown in Example 1.9, there is no essential difference between $\mathbb{H}^{3}$ and the hyperbolic $3$-space with negative constant sectional curvature not equal to 1-1. sectional curvature 1-1 and the related volume element $dv$. For a fixed point $x_{0}\in\mathbb{H}^{3}$, set $r:=d_{\mathbb{H}^{3}}(x_{0},y)$, which measures the hyperbolic distance between two points $x_{0}$, $y$ in $\mathbb{H}^{3}$. If the potential function $\phi$ was chosen to be
@@
Readers can check [36, Subsection 4.6] for the information on the weighted heat kernel $H^{\phi}$ in hyperbolic spaces with other dimensions.
@@
That is to say, under what kind of assumptions, one has $l=\infty$. This question is important because if $l=\infty$, the weighted heat kernel comparison (1.7) or (1.8) might sustain continuously without any interruption. By the Sturm-Picone comparison theorem, it is easy to know that if $\kappa(s)\leq 0$, then (1.5) has a unique solution on $(0,\infty)$, i.e. $l=\infty$. What about other situations? In [11], [46, Chapter 2, Section 2.6], some sufficient conditions involving $\kappa(s)$ have been given such that the initial value problem (1.5) has a unique solution on $(0,\infty)$. To the best of our knowledge, so for, the sufficient and necessary condition of the result $l=\infty$ for the initial value problem (1.5) has not been given yet.
@@
If $\phi=const.$, then the weighted heat kernel $H^{\phi}(\cdot,\cdot,\cdot)$ becomes the usual heat kernel $H(\cdot,\cdot,\cdot)$ of the heat equation, and correspondingly, Theorem 1.6 degenerates into our previous work [47, Theorem 6.6] exactly.
@@
In [47, Theorem 6.6], the author has given lower and upper bounds for the heat kernel of geodesic balls on complete manifolds with radial (Ricci and sectional) curvatures bounded. This conclusion extends Cheeger-Yau’s lower bound estimate [18] and Debiard-Gaveau-Mazet’s upper bound estimate [27], respectively. Hence, one can easily see from Corollary 1.14 that our Theorem 1.6 here improves the heat kernel estimates in [18, 27, 47] a lot.
@@
As mentioned in [47, Remark 6.7], the completeness of the prescribed manifold $M^{n}$ is a little strong to get comparison results for the heat kernel, since Cheeger-Yau [18] have shown that if the injectivity radius at some point $q$ of a prescribed manifold $M^{n}$ is bounded from below, then under the assumptions on curvature therein, a lower bound can be given for the heat kernel of geodesic balls on $M^{n}$. Maybe the completeness of $M^{n}$ is also a little strong to get comparisons (1.7) and (1.8) for the weighted heat kernel. However, here we wish to assume that $M^{n}$ is complete, this is because that if $M^{n}$ is complete, then for $B(q,r_{0})\subseteq M^{n}$ with $r_{0}$ finite we can always find optimally continuous bounds for the radial Ricci and sectional curvatures with respect to $q$ (see (2.14) and (2.15)). Therefore, the assumption on the completeness of $M^{n}$ is feasible.
@@
with $\partial\Omega$ the boundary777 In general case, to avoid a little bit boring discussion on the regularity of the boundary $\partial\Omega$, which is obviously not the point mainly considered in this paper, one might assume that $\partial\Omega$ is smooth for the Dirichlet eigenvalue problem of the Witten-Laplacian. However, if one checks carefully the details in this paper, he or she would find that weaker regularity assumptions on $\partial\Omega$, such as $\partial\Omega$ is piecewise smooth or Lipschitz continuous, can also make sure that the spectral results of this paper are valid. Based on this reason, in this paper, we do not make any explicit assumption for the regularity of $\partial\Omega$ and we hope that one could understand that a suitable regularity assumption for $\partial\Omega$ has been used potentially in our paper. of the bounded domain $\Omega$. For this problem, one has (see e.g. [20, pp. 1247-1248]):
@@
In our previous work [21, Introduction], we have given a detailed explanation on why is interesting and important to study spectral geometric problems related to the Witten-Laplacian. We already have some interesting works about spectral isoperimetric inequalities, spectral estimates and geometric functional inequalities related to the Witten-Laplacian – see, e.g., [20, 21, 28, 48, 50, 68].
@@
As applications of the weighted heat kernel comparison (i.e. Theorem 1.6), we can get the following eigenvalue comparison theorems of the Witten-Laplacian, which can be seen as the weighted version of those shown in [31, Theorems 3.6 and 4.4].
@@
respectively. Moreover, when the equalities in the above two eigenvalue inequalities hold, the corresponding isometric rigidity between two geodesic balls can also be attained. This conclusion is exactly the assertions in [31, Theorems 3.6 and 4.4].
@@
In [59, Section 4], by mainly using a similar argument to that in Cheng’s proof of [23, Theorem 1.1], A. G. Setti obtained an eigenvalue comparison theorem for the first Dirichlet eigenvalue of the Witten-Laplacian on geodesic balls of complete Riemannian $n$-manifolds with a modified Ricci curvature bounded from below by a constant. More precisely, for a complete Riemannian $n$-manifold $M^{n}$ with the metric $\langle\cdot,\cdot\rangle$, let $w$ be a given smooth strictly positive function on $M^{n}$ (which was called weight function therein), and then A. G. Setti defined a modified Ricci curvature as follows
@@
Therefore, it follows from (1.21) that comparing with the Bakry–Émery Ricci curvature tensor $\mathrm{Ric}_{\phi}$, the modified Ricci curvature tensor $\mathrm{Ric}_{w}$ defined in [59] has an extra term $d\phi\otimes d\phi$. Besides, the only difference between $\mathrm{Ric}_{w}$ and the $N$-dimensional Bakry-Émery Ricci curvature tensor $\mathrm{Ric}_{\phi}^{N}$ is the coefficient $\frac{1}{N-n}$ in front of the term $d\phi\otimes d\phi$, which in my opinion is not an essential difference. In [59, Theorem 4.2], A. G. Setti showed that if
@@
with the weight function $w$ satisfying $w=e^{-\phi}$. That is to say, the first Dirichlet eigenvalue $\lambda_{1,\phi}(B(q,R))$ of the Witten-Laplacian $\Delta_{\phi}$ on the geodesic ball $B(q,R)\subseteq M^{n}$ is bounded from above by the first Dirichlet eigenvalue $\lambda_{1}\left(\mathcal{B}_{n+1}(\alpha/n,R)\right)$ of the Laplacian on a geodesic ball $\mathcal{B}_{n+1}(\alpha/n,R)$ in the $(n+1)$-dimensional space form $\mathbb{M}^{n+1}(\alpha/n)$. Specially, in [59, Section 4], he also mentioned that if furthermore $\alpha>0$, then using a Bishop-type volume comparison theorem [59, Theorem 4.1], one knows that $d_{M^{n}}(q,\hat{q})\leq\pi\sqrt{\frac{n}{\alpha}}$, $\forall\hat{q}\in M^{n}$. Hence, if $R>\pi\sqrt{\frac{n}{\alpha}}$, then $B(q,R)=M^{n}$ and $\mathcal{B}_{n+1}(\alpha/n,R)=\mathbb{M}^{n+1}(\alpha/n)$, which implies $\lambda_{1,\phi}(B(q,R))=0=\lambda_{1}\left(\mathcal{B}_{n+1}(\alpha/n,R)\right)$, and the comparison (1.22) holds trivially. That is to say, the eigenvalue comparison (1.22) does not have the rigidity characterization for the equality “==”, however in some special cases the equality can be achieved.

### v2

# Weighted heat kernel comparison theorems and its applications in spectral geometry

@@ 1 Introduction
In the past four decades, smooth metric measure spaces have been extensively investigated in geometric analysis, and have close relation with probability theory and optimal transport. We wish to mention again several important examples listed in our previous work [72, Section 1] to support this viewpoint — G. Perelman’s pioneering work [60] on the entropy formula of the Ricci flow under a weighted measure, Munteanu-Wang’s series works [57, 58, 59] about geometry of manifolds with weighted densities, Wei-Wylie’s interesting work [68] on some comparison theorems for manifolds with Bakry-Émery Ricci curvature bounded. Besides, recently, by mainly using the rearrangement technique and suitably constructing trial functions, under the constraint of fixed weighted volume, R. F. Chen and J. Mao [21] successfully obtain several isoperimetric inequalities for the first and the second Dirichlet eigenvalues, the first nonzero Neumann eigenvalue of the Witten Laplacian on bounded domains in space forms. These spectral isoperimetric inequalities extend the classical ones (i.e., the Faber-Krahn inequality, the Hong-Krahn-Szegő inequality, and the Szegő-Weinberger inequality) of the Laplacian.
@@
$\lim\limits_{t\rightarrow 0}\int\limits_{M^{n}}f(x)H^{\phi}(x,y,t)e^{-\phi}dv(x)=f(y).$
@@
For detailed calculations, see [70, Appendix].
@@
There are so many mathematicians having interest on the (non-weighted or weighted) heat kernel and its applications, and many interesting results have been obtained. Here we wish to give a partial list of those results which influence me a lot. In 1981, S. Y. Cheng, P. Li and S. T. Yau [26] gave uniform Gaussian estimates for the heat kernel on Riemannian manifolds with sectional curvature bounded from below. Their estimates were improved by J. Cheeger, M. Gromov, M. Taylor [17] to manifolds with bounded geometry. By using the gradient estimate technique and the Harnack inequality, P. Li and S. T. Yau [44] obtained sharp Gaussian upper and lower bounds for the Schrödinger operator on Riemannian manifolds with nonnegative Ricci curvature. For smooth metric measure spaces with $\mathrm{Ric}^{N}_{\phi}$ ($N<\infty$) bounded from below by a negative quadratic function, X. D. Li [43] got Gaussian estimates for the weighted heat kernel, and proved an $L^{1}_{\phi}$-Liouville theorem. In [69], J. Y. Wu obtained weighted heat kernel estimates on smooth metric measure spaces with the weighted Ricci curvature bounded from below by a negative constant and the potential function bounded. By using the De Giorgi-Nash-Moser theory and the weighted version of Davies’s integral estimate [27], J. Y. Wu and P. Wu [70, Theorem 1.1] successfully gave a local Gaussian upper bound for the weighted heat kernel on complete noncompact smooth metric measure spaces $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$ with $\mathrm{Ric}_{\phi}\geq 0$. Their upper bound estimate generalizes the classical result of Li-Yau [44]. As applications of their local Gaussian upper bound estimate for the weighted heat kernel, they showed that if furthermore the potential function $\phi$ has at most quadratic growth, then a sharp $L^{1}_{\phi}$-Liouville theorem for $\phi$-subharmonic functions and an $L^{1}_{\phi}$-uniqueness property for nonnegative solutions of the weighted heat equation can be obtained (see [70, Theorem 1.5]). Besides, it is not hard to see from [14, 40] that any complete noncompact shrinking or steady gradient Ricci soliton has nonnegative weighted Ricci curvature and its potential function $\phi$ should have at most quadratic growth. Then together with Wu-Wu’s $L^{1}_{\phi}$-Liouville theorem mentioned above, it is easy to know that any $L^{1}_{\phi}$-integral $\phi$-subharmonic function on gradient shrinking and steady Ricci solitons must be constant (see [70, Corollary 1.6]). Later, as a sequel of the work [70], by mainly using the De Giorgi-Nash-Moser theory, J. Y. Wu and P. Wu [71, Theorem 1.1] obtained local Gaussian (upper and lower) bounds for the weighted heat kernel on complete noncompact smooth metric measure spaces $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$ with Bakry-Émery Ricci curvature bounded from below by some nonnegative constant (i.e. $\mathrm{Ric}_{\phi}\geq-(n-1)K$ for some constant $K\geq 0$). Moreover, when the curvature assumption was reduced to $\mathrm{Ric}_{\phi}\geq 0$ (i.e. $K=0$), they pointed out that the local Gaussian lower bound for the weighted heat kernel can be achieved by the $\phi$-heat kernel for steady Gaussian soliton (explicitly given in Example 1.3) as long as the time $t$ is large enough. Besides, as applications, they used their local Gaussian (upper and lower) bounds for the weighted heat kernel to obtain estimates for closed eigenvalues and for $\phi$-Green’s function of the Witten-Laplacian – see [71, Sections 6 and 7] for details. In our recent work [72], by utilizing some functional inequalities and combining with De Giorgi-Moser iteration technique, some estimates for the weighted $L^{\infty}$-norm of eigenfunctions of the Witten-Laplacian on smooth metric measure spaces have been given. Besides, we also proved the regularity and uniqueness of the weighted heat kernel on smooth metric measure spaces, which generalized Chavel-Feldman’s related results in [15].
@@
In this paper, we wish to give a comparison theorem for the weighted heat kernel (of the weighted heat equation) on complete manifolds with radial curvatures bounded. In this setting, the model space would be spherically symmetric manifolds. In order to state our main result clearly, we need to recall some preliminary knowledge about the exponential mapping, the cut locus, the radial direction, the injectivity radius, the radial (Ricci or sectional) curvature having a bound given by a continuous function (see Definitions 2.2 and 2.3), spherically symmetric manifolds (see Definition 2.1), and so on. However, in order to control the length of this section, we wish to give preliminaries in Section 2 of this paper. By the way, these preliminaries can also be found in our previous works [33, Section 1], [48, Section 2.1 of Chapter 2], [49, Section 2].
@@
which will be used to determine our model space $0,l)\times_{f}\mathbb{S}^{n-1}$ (i.e. spherically symmetric manifolds, which are called generalized space forms in [42]). See Definition 2.1 in Section 2 for the precise statement such that a domain to be spherically symmetric. Readers can check our previous work [33, pp. 705–709] for a brief introduction and an interesting spectral asymptotic property of spherically symmetric manifolds. Obviously, if in (1.5), $\kappa(s)\equiv K$ for some constant $K$, then
@@
(see [38, Subsection 4.6]) Let $\mathbb{H}^{3}$ be the hyperbolic $3$-space with constant666 For the weighted heat kernel $H^{\phi}$ shown in Example 1.9, there is no essential difference between $\mathbb{H}^{3}$ and the hyperbolic $3$-space with negative constant sectional curvature not equal to 1-1. sectional curvature 1-1 and the related volume element $dv$. For a fixed point $x_{0}\in\mathbb{H}^{3}$, set $r:=d_{\mathbb{H}^{3}}(x_{0},y)$, which measures the hyperbolic distance between two points $x_{0}$, $y$ in $\mathbb{H}^{3}$. If the potential function $\phi$ was chosen to be
@@
Readers can check [38, Subsection 4.6] for the information on the weighted heat kernel $H^{\phi}$ in hyperbolic spaces with other dimensions.
@@
That is to say, under what kind of assumptions, one has $l=\infty$. This question is important because if $l=\infty$, the weighted heat kernel comparison (1.7) or (1.8) might sustain continuously without any interruption. By the Sturm-Picone comparison theorem, it is easy to know that if $\kappa(s)\leq 0$, then (1.5) has a unique solution on $(0,\infty)$, i.e. $l=\infty$. What about other situations? In [12], [48, Chapter 2, Section 2.6], some sufficient conditions involving $\kappa(s)$ have been given such that the initial value problem (1.5) has a unique solution on $(0,\infty)$. To the best of our knowledge, so for, the sufficient and necessary condition of the result $l=\infty$ for the initial value problem (1.5) has not been given yet.
@@
If $\phi=const.$, then the weighted heat kernel $H^{\phi}(\cdot,\cdot,\cdot)$ becomes the usual heat kernel $H(\cdot,\cdot,\cdot)$ of the heat equation, and correspondingly, Theorem 1.6 degenerates into our previous work [49, Theorem 6.6] exactly.
@@
In [49, Theorem 6.6], the author has given lower and upper bounds for the heat kernel of geodesic balls on complete manifolds with radial (Ricci and sectional) curvatures bounded. This conclusion extends Cheeger-Yau’s lower bound estimate [19] and Debiard-Gaveau-Mazet’s upper bound estimate [28], respectively. Hence, one can easily see from Corollary 1.14 that our Theorem 1.6 here improves the heat kernel estimates in [19, 28, 49] a lot.
@@
As mentioned in [49, Remark 6.7], the completeness of the prescribed manifold $M^{n}$ is a little strong to get comparison results for the heat kernel, since Cheeger-Yau [19] have shown that if the injectivity radius at some point $q$ of a prescribed manifold $M^{n}$ is bounded from below, then under the assumptions on curvature therein, a lower bound can be given for the heat kernel of geodesic balls on $M^{n}$. Maybe the completeness of $M^{n}$ is also a little strong to get comparisons (1.7) and (1.8) for the weighted heat kernel. However, here we wish to assume that $M^{n}$ is complete, this is because that if $M^{n}$ is complete, then for $B(q,r_{0})\subseteq M^{n}$ with $r_{0}$ finite we can always find optimally continuous bounds for the radial Ricci and sectional curvatures with respect to $q$ (see (2.14) and (2.15)). Therefore, the assumption on the completeness of $M^{n}$ is feasible.
@@
with $\partial\Omega$ the boundary777 In general case, to avoid a little bit boring discussion on the regularity of the boundary $\partial\Omega$, which is obviously not the point mainly considered in this paper, one might assume that $\partial\Omega$ is smooth for the Dirichlet eigenvalue problem of the Witten-Laplacian. However, if one checks carefully the details in this paper, he or she would find that weaker regularity assumptions on $\partial\Omega$, such as $\partial\Omega$ is piecewise smooth or Lipschitz continuous, can also make sure that the spectral results of this paper are valid. Based on this reason, in this paper, we do not make any explicit assumption for the regularity of $\partial\Omega$ and we hope that one could understand that a suitable regularity assumption for $\partial\Omega$ has been used potentially in our paper. of the bounded domain $\Omega$. For this problem, one has (see e.g. [21, pp. 1247-1248]):
@@
In our previous work [22, Introduction], we have given a detailed explanation on why is interesting and important to study spectral geometric problems related to the Witten-Laplacian. We already have some interesting works about spectral isoperimetric inequalities, spectral estimates and geometric functional inequalities related to the Witten-Laplacian – see, e.g., [21, 22, 30, 50, 52, 73].
@@
As applications of the weighted heat kernel comparison (i.e. Theorem 1.6), we can get the following eigenvalue comparison theorems of the Witten-Laplacian, which can be seen as the weighted version of those shown in [33, Theorems 3.6 and 4.4].
@@
respectively. Moreover, when the equalities in the above two eigenvalue inequalities hold, the corresponding isometric rigidity between two geodesic balls can also be attained. This conclusion is exactly the assertions in [33, Theorems 3.6 and 4.4].
@@
In [63, Section 4], by mainly using a similar argument to that in Cheng’s proof of [24, Theorem 1.1], A. G. Setti obtained an eigenvalue comparison theorem for the first Dirichlet eigenvalue of the Witten-Laplacian on geodesic balls of complete Riemannian $n$-manifolds with a modified Ricci curvature bounded from below by a constant. More precisely, for a complete Riemannian $n$-manifold $M^{n}$ with the metric $\langle\cdot,\cdot\rangle$, let $w$ be a given smooth strictly positive function on $M^{n}$ (which was called weight function therein), and then A. G. Setti defined a modified Ricci curvature as follows
@@
Therefore, it follows from (1.21) that comparing with the Bakry–Émery Ricci curvature tensor $\mathrm{Ric}_{\phi}$, the modified Ricci curvature tensor $\mathrm{Ric}_{w}$ defined in [63] has an extra term $d\phi\otimes d\phi$. Besides, the only difference between $\mathrm{Ric}_{w}$ and the $N$-dimensional Bakry-Émery Ricci curvature tensor $\mathrm{Ric}_{\phi}^{N}$ is the coefficient $\frac{1}{N-n}$ in front of the term $d\phi\otimes d\phi$, which in my opinion is not an essential difference. In [63, Theorem 4.2], A. G. Setti showed that if
@@
with the weight function $w$ satisfying $w=e^{-\phi}$. That is to say, the first Dirichlet eigenvalue $\lambda_{1,\phi}(B(q,R))$ of the Witten-Laplacian $\Delta_{\phi}$ on the geodesic ball $B(q,R)\subseteq M^{n}$ is bounded from above by the first Dirichlet eigenvalue $\lambda_{1}\left(\mathcal{B}_{n+1}(\alpha/n,R)\right)$ of the Laplacian on a geodesic ball $\mathcal{B}_{n+1}(\alpha/n,R)$ in the $(n+1)$-dimensional space form $\mathbb{M}^{n+1}(\alpha/n)$. Specially, in [63, Section 4], he also mentioned that if furthermore $\alpha>0$, then using a Bishop-type volume comparison theorem [63, Theorem 4.1], one knows that $d_{M^{n}}(q,\hat{q})\leq\pi\sqrt{\frac{n}{\alpha}}$, $\forall\hat{q}\in M^{n}$. Hence, if $R>\pi\sqrt{\frac{n}{\alpha}}$, then $B(q,R)=M^{n}$ and $\mathcal{B}_{n+1}(\alpha/n,R)=\mathbb{M}^{n+1}(\alpha/n)$, which implies $\lambda_{1,\phi}(B(q,R))=0=\lambda_{1}\left(\mathcal{B}_{n+1}(\alpha/n,R)\right)$, and the comparison (1.22) holds trivially. That is to say, the eigenvalue comparison (1.22) does not have the rigidity characterization for the equality “==”, however in some special cases the equality can be achieved.
@@
On $1^{\mathrm{st}}$ March 2026, we posted the first version of this paper on arXiv. However, on $12^{\mathrm{th}}$ March 2026, we noticed the work [64] which was just posted on arXiv on $9^{\mathrm{th}}$ March 2026 (i.e. almost three days ago), and has cited my two works [29, 49]. We found that the assertion in Theorem 1.18 can be improved to the case of weighted $p$-Laplacian by making some adjustments to the argument in the proof of [64, Theorem 1.4]. We wish to write this content as an Appendix C in the second version of this paper.